# Graph-based Approaches and Functionalities in Retrieval-Augmented Generation: A Comprehensive Survey

### v1

@@ 8. Future Research Opportunities
Incorporating User Interaction. Future work in RAG systems can be significantly improved by integrating user interaction to improve the efficiency, accuracy, and user satisfaction of knowledge retrieval. By integrating Human-Computer Interaction (HCI) principles, RAG systems can offer more intuitive and adaptive interfaces, allowing users to actively shape the retrieval process. For example, adaptive query refinement (deshpande2007adaptive, ) could enable users to iteratively clarify or expand their queries, with the system responding to these modifications in real-time. Additionally, interactive visualizations of the knowledge graph could allow users to explore relationships between data points, enhancing understanding and control over the retrieval process. Incorporating user feedback loops (mansoury2020feedback, ) — where users can rate the relevance of results or suggest corrections—would also help the system learn and adapt over time, making it more responsive to evolving user needs. Moreover, adopting context-aware systems (hong2009context, ) could allow RAG systems to interpret not only the user’s query but also the broader context of the interaction, further refining results based on user preferences and past behaviors. By blending these HCI concepts into RAG design, future systems can offer more personalized, efficient, and transparent experiences, ultimately enhancing user trust and satisfaction.
Robustness and Bias Mitigation. In Graph-based RAG systems, ensuring robustness and bias mitigation presents unique challenges, not only because of the underlying hallucination and unreliability issues in LLMs but also due to the graph structure itself. While graphs provide a more organized and transparent representation of knowledge, they can also introduce their own biases and vulnerabilities (tang2020unbiased, ). For instance, when the graph’s structure reflects biased relationships or underrepresents certain concepts, this bias will be directly transferred into the retrieval process, leading to skewed or unfair results. Furthermore, the graph’s robustness can be compromised if it includes faulty or incomplete data, which could cause incorrect retrievals or unreliable outputs. To address these challenges, future work can focus on developing bias-aware graph construction techniques, such as ensuring diverse and balanced knowledge representation within the graph’s nodes and edges (dong2022edits, ). Additionally, robustness-enhancing strategies like graph regularization (zhu2021shift, ) can help improve the system’s resilience by addressing missing or corrupted data.
Practical Applications in Industry. Currently, graph-based RAG methods are widely studied in research but have yet to see widespread adoption in industry. The key barriers to industrial deployment include challenges related to efficiency, accuracy, and scalability. For instance, while the graph structure offers advantages in organizing and retrieving knowledge, it can also become computationally expensive and difficult to scale in real-world applications with large, dynamic datasets. However, initial industry efforts have demonstrated promising results: Microsoft’s GraphRAG  [^3] https://github.com/microsoft/graphrag successfully leverages enterprise knowledge graphs to improve internal question-answering accuracy; Writer  [^4] https://writer.com/ utilizes graph-based RAG to enhance enterprise content generation and Q&A systems; retailers like Leroy Merlin  [^5] https://www.lettria.com/case-study/leroy-merlin-knowledge-graph-product-recommendations use product knowledge graphs to refine recommendation systems, providing personalized customer experiences. To further bridge the gap between research and practical industry use, future work should focus on optimizing graph-based RAG systems for real-time performance and large-scale data handling, ensuring that they can handle vast amounts of data quickly and accurately. Additionally, adapting these systems to industry-specific domains and use cases will be essential, as the needs in healthcare, finance, and e-commerce, for example, may require tailored retrieval mechanisms and specialized graph structures. By addressing these challenges and optimizing the system for industrial environments, future research can make graph-based RAG systems more accessible and applicable to a broader range of industries, driving the practical adoption of these advanced knowledge retrieval methods.

### v2

@@ 10. Future Research Opportunities
Incorporating User Interaction. Future work in RAG systems can be significantly improved by integrating user interaction to improve the efficiency, accuracy, and user satisfaction of knowledge retrieval. By integrating Human-Computer Interaction (HCI) principles, RAG systems can offer more intuitive and adaptive interfaces, allowing users to actively shape the retrieval process. For example, adaptive query refinement (Deshpande et al., 2007) could enable users to iteratively clarify or expand their queries, with the system responding to these modifications in real-time. Additionally, interactive visualizations of the knowledge graph could allow users to explore relationships between data points, enhancing understanding and control over the retrieval process. Incorporating user feedback loops (Mansoury et al., 2020) — where users can rate the relevance of results or suggest corrections—would also help the system learn and adapt over time, making it more responsive to evolving user needs. Moreover, adopting context-aware systems (Hong et al., 2009) could allow RAG systems to interpret not only the user’s query but also the broader context of the interaction, further refining results based on user preferences and past behaviors. By blending these HCI concepts into RAG design, future systems can offer more personalized, efficient, and transparent experiences, ultimately enhancing user trust and satisfaction.